Please help me solve this equation the sum of two consecutive integers is 19 less than their product. find the integers.
Answer by ewatrrr(24785)


Hi
two consecutive integers
Let x and (x+1) represent the two consecutive integers
Question states***
x +(x+1)= x(x+1)-19
Solving for x
2x + 1 = x^2 + x - 19
x^2 -x - 20 = 0
factoring
(x-5)(x+4)= 0 Note:SUM of the inner product(-5x) and the outer product(4x) = -x
(x-5)=0 x = 5, the two consecutive integers are 5 and 6
(x+4)= 0 x = -4 the two consecutive integers are -4 and -3
CHECKING our Answer(s)***
5 + 6 = 30 - 19
-4 - 3 = -4*-3 - 19